Write a system of equations to solve the following problem. Let c be the number of child tickets and a be the number of adult tickets.
Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $150, and 40 tickets were sold, how many of each type of ticket was sold?

Answer by ankor@dixie-net.com(22740)

Write a system of equations to solve the following problem.
Let c be the number of child tickets and a be the number of adult tickets.
Each child ticket for a ride costs $3, while each adult ticket costs $5.
If the ride collected a total of $150, and 40 tickets were sold, how many
of each type of ticket was sold?
:
Two equations:
c + a = 40
a = (40-c); use this for substitution
and
3c + 5a = $150